# A Skeleton-Based Topological Planner for Exploration in Complex Unknown Environments

### v2

@@ III-B Incremental Skeleton Extraction
However, extracting a strict GVD as per the definition incurs high computational complexity, and such precision is unnecessary for exploration tasks. Inspired by [23], we propose an approximate method to iteratively identify skeleton candidates and support incremental updates. Pseudo-code for incremental skeleton extraction is provided in Alg. 1. The 8-neighborhoods of $s(i,j)$, $Adj_{8}(s)$, consist of the points surrounding $s(i,j)$, with $n_{1}$ as the top neighbor and $n_{2}$ to $n_{8}$ indexed clockwise.
@@ 6       $changed\leftarrow\text{false}$;
9                   if $waveFront_{s}>iteration$ or ${state(Skeleton,s)\neq\text{free}}$ then
@@
11
12                  if ($phase=1$ and $(a)\land(b)\land(c)\land(d)$) or
13                   ($phase=2$ and $(a)\land(b)\land(e)\land(f)$) then
@@ 26             $phase\leftarrow 3-phase$;
To support incremental updates, the wavefront of each removed grid is recorded, and its state is marked based on the states of its neighboring grids (lines 13-18, Alg. 1). During each iteration, grids with wavefront values exceeding the current iteration count are skipped (line 9, Alg. 1). Upon completion, $waveFront$ values of unknown grids are reset to zero (line 22, Alg. 1). This approach eliminates redundant wavefront propagation from obstacles while addressing unknown region uncertainty, ensuring computational efficiency and skeleton continuity, as shown in Fig. 2. Additionally, the wavefront value provides insight into the path’s width and is subsequently applied in Sec. III-C2 and Sec. III-D.
@@ V CONCLUSION
In this paper, we introduced STGPlanner, a novel autonomous exploration framework based on skeletal topological graph (STG). We proposed an incremental skeleton extraction method that allows the STG to expand effectively. Further, a FSM was designed to encapsulate the entire exploration strategy. By leveraging the topological structure of the environment, our method significantly reduces BFMs, enhancing exploration efficiency while minimizing computational costs. However, several limitations remain. In larger open environments, skeletal extraction may prove ineffective, necessitating a reliance on traditional frontier-based guidance methods. Furthermore, sensor noise and dynamic elements in real-world scenarios can adversely affect the skeletal structure. Future work will focus on adapting the system for dynamic environments and facilitating multi-robot cooperation.

### v3

@@ III-B Incremental Skeleton Extraction
However, extracting a strict GVD as per the definition incurs high computational complexity, and such precision is unnecessary for exploration. Inspired by [23], we propose an approximate method to iteratively identify skeleton and support incremental updates. Pseudo-code for incremental skeleton extraction is provided in Alg. 1. The 8-neighborhoods of $s(i,j)$, $Adj_{8}(s)$, consist of the points surrounding $s(i,j)$, with $n_{1}$ as the top neighbor and $n_{2}$ to $n_{8}$ indexed clockwise.
@@ 6       $changed\leftarrow\text{false}$;
8             $condition=(Skeleton=\mathrm{free})$ and $waveFront\geq iteration$;
@@
10                   if ${state(Skeleton,s)\neq\text{free}}$ then
@@
12
13                  if ($phase=1$ and $(a)\land(b)\land(c)\land(d)$) or ($phase=2$ and $(a)\land(b)\land(e)\land(f)$) then
@@ 26             $phase\leftarrow 3-phase$;
To support incremental updates, each grid cell that satisfies the specified condition not only updates its state but also records the current iteration as its wavefront. This value is then used to determine the parameter $n_{i}$ in Eq. (3) via $n_{i}=condition_{s_{n_{i}}}$ (lines 8-13, Alg. 1). Moreover, unknown regions are treated as special occupied grids; all grids covered by waves triggered from these grids are marked as unknown, and, ultimately, $waveFront$ and $Skeleton$ within these unknown regions are reset (lines 15-18, 22, Alg. 1). This procedure both avoids redundant computation of obstacle-triggered wavefront propagation and accounts for the uncertainty of unknown areas, thereby ensuring the coherence and stability of skeleton, as shown in Fig. 2. Additionally, $waveFront$ provides insight into the path’s width and is subsequently applied in Secs. III-C2 and III-D.
@@ V CONCLUSION
In this paper, we introduced STGPlanner, a novel autonomous exploration framework based on skeletal topological graph (STG). We proposed an incremental skeleton extraction method that allows the STG to expand effectively. Further, a FSM was designed to encapsulate the entire exploration strategy. By leveraging the topological structure of the environment, our method significantly reduces BFMs, enhancing exploration efficiency while minimizing computational costs. However, several limitations remain. In larger open environments, skeletal extraction may prove ineffective, necessitating a reliance on traditional frontier-based methods. Furthermore, sensor noise and dynamic elements in real-world scenarios can adversely affect the skeletal structure. Future work will focus on adapting the system for dynamic environments and facilitating multi-robot cooperation.